# Determination of the stably free cancellation property for orders

### v2

@@ 3.1. Stably free and locally free cancellation
$X\oplus\Lambda^{(m)}\cong\Lambda^{(m+n)}\textrm{ for some }m,n\geq 0\implies X\cong\Lambda^{(n)}.$ (3.1)
@@
$X\oplus\Lambda^{(m)}\cong Y\oplus\Lambda^{(m)}\textrm{ for some }m\geq 0\implies X\cong Y.$ (3.2)
@@
If $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(m)}\cong\Lambda^{(m+n)}\textrm{ for some }m,n\geq 0$ then it is straightforward to see that $X$ is a $\Lambda$-lattice; since $X$ is stably free it is also locally free (see §2.4). Therefore if $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders that have SFC but do not have LFC (see [Sme15] or [SV19]).
By (2.2), $\Lambda$ has SFC if and only if (3.1) holds for all $X\in g(\Lambda)$. Similarly, $\Lambda$ has LFC if and only if (3.2) holds for all $X,Y\in g(\Lambda)$. Hence the canonical surjective map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is bijective precisely when $\Lambda$ has LFC. Moreover, $\Lambda$ has SFC precisely when the fibre above 00 has exactly one element, that is, when every stably free $\Lambda$-lattice of rank $1$ is in fact free (of rank $1$).
@@
The claim for LFC was proven by Fröhlich [Frö75, VIII], whose approach was subsequently refined by Swan. Here we adapt the proof of [Nic21a, Theorem 1.1], which shows the result for SFC in the case that $\Lambda$ and $\Gamma$ are integral group rings. By [Swa83, Theorem A10] the diagram
@@
is a weak pull back in the sense that a compatible pair of elements in $\mathrm{Cl}(\Lambda)$ and $\mathrm{LF}_{1}(\Gamma)$ has a not-necessarily-unique lift to $\mathrm{LF}_{1}(\Lambda)$. It follows that the fibre in $\mathrm{LF}_{1}(\Lambda)$ over an element $x\in\mathrm{Cl}(\Lambda)$ maps onto the fibre in $\mathrm{LF}_{1}(\Gamma)$ over the image of $x$ in $\mathrm{Cl}(\Gamma)$. Thus if $\Lambda$ has LFC then $\Gamma$ has LFC. Taking $x=0$ shows that if $\Lambda$ has SFC then $\Gamma$ has SFC. ∎
@@
Suppose that $\Lambda=\Lambda_{1}\oplus\Lambda_{2}$ is a direct sum of two $R$-orders $\Lambda_{1}$ and $\Lambda_{2}$. Then $\Lambda$ has SFC (resp. LFC) if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC (resp. LFC).
@@ 3.2. Totally definite quaternion algebras and the Eichler condition
Let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. A totally definite quaternion algebra is a simple finite-dimensional $K$-algebra $B$ whose completion at every archimedean prime of the centre of $B$ is isomorphic to $\mathbb{H}$ (in particular, this forces $K$ to be totally real).
We say that a finite-dimensional semisimple $K$-algebra $A$ satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra. A proof of the following result can be found in [CR87, (51.24)] or [Jac68, §4].
@@
In [SV19] it was shown that the question of whether a given order in a totally definite quaternion algebra has SFC or LFC can be tackled algorithmically.
@@
There exists an algorithm that given an $R$-order $\Lambda$ in a totally definite quaternion algebra over $K$, determines whether $\Lambda$ has SFC (resp. LFC).
@@
There exists an algorithm that given a maximal $R$-order $\Lambda$ in a finite-dimensional semisimple $K$-algebra, determines whether $\Lambda$ has SFC (resp. LFC).
@@
Let $\mathcal{M}$ be such a maximal $R$-order. Then we can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $R$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. Therefore the desired result now follows from Corollary 3.3, Theorem 3.4 and Theorem 3.5. ∎
@@ 4. Review of cancellation properties for integral group rings
We recall results concerning SFC and LFC for group rings $\mathbb{Z}[G]$ of a finite group $G$.
@@
The finite non-cyclic subgroups of $\mathbb{H}^{\times}$ are called the binary polyhedral groups. These are the generalised quaternion groups $Q_{4n}$ for $n\geq 2$ or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups (see [CR87, §51A] and the references therein). It is well known that a finite group $G$ has a binary polyhedral quotient precisely when the group algebra $\mathbb{Q}[G]$ does not satisfy the Eichler condition (see [Nic21a, Proposition 1.5] or [CR87, (51.3)], for example).
@@
$Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,Q_{24},\,Q_{28},\,Q_{36},\,Q_{60},\,\tilde{T},\,\tilde{O},\,\tilde{I}.$ (4.1)
@@
For a finite group $G$ and a field $\mathbb{F}$, let $r_{\mathbb{F}}(G)$ denote the number of irreducible $\mathbb{F}$-representations of $G$, and let $SK_{1}(\mathbb{Z}[G])=\ker(K_{1}(\mathbb{Z}[G])\rightarrow K_{1}(\mathbb{Q}[G]))$; this is a finite abelian group by a theorem of Bass [CR87, (45.20)]. We say that $K_{1}(\mathbb{Z}[G])$ is represented by trivial units if the canonical map ±$\pm G\rightarrow K_{1}(\mathbb{Z}[G])$ is surjective.
@@
5. Fibre products and reduction criteria for SFC
@@
5.1. Fibre products
Suppose that there is a fibre product
${\Lambda}$${\Lambda_{1}}$${\Lambda_{2}}$${\overline{\Lambda}.}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (5.1)
@@ 5.1. Fibre products
$\Lambda\cong\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\quad\textrm{ and }\quad A\cong A_{1}\oplus A_{2},$ (5.2)
@@
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\overline{\Lambda}$ is $R$-torsion and there is a fibre product as in (5.1) such that each of $\pi_{1},\pi_{2},g_{1},g_{2}$ is the relevant canonical surjection.
@@ 5.2. Review of results of Reiner–Ullom
Given a fibre product as in (5.1), assume that at least one of $g_{1}$ and $g_{2}$ is surjective. For $u\in\overline{\Lambda}^{\times}$, define
@@ 5.3. A reduction criterion for SFC
The following underpins many of our results.
@@
Assume the setting of Corollary 5.3. Then $\Lambda$ has SFC if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and $W=\ker(\partial)$.
@@
Suppose conversely that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W=\ker(\partial)$. Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free of rank $1$. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $[X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. Then for $i=1,2$ we have $[\Lambda_{i}\otimes_{\Lambda}X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda_{i})$ and so $\Lambda_{i}\otimes_{\Lambda}X\cong\Lambda_{i}$ since $\Lambda_{i}$ has SFC. Hence $X\cong M(u)$ for some $u\in\overline{\Lambda}^{\times}$ by Proposition 5.2 (iv). Since $X$ is stably free, Corollary 5.3 (iii) implies that $u\in\ker(\partial)$. Thus since $W=\ker(\partial)$, Corollary 5.3 (ii) implies that $M(u)\cong\Lambda$. Therefore $X\cong M(u)\cong\Lambda$. ∎
@@
Assume the setting of Corollary 5.3. Suppose that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W$ is a group. Let $N$ be a normal subgroup of $\ker(\partial)$ contained in $W$ and let $S$ be a generating subset of $\ker(\partial)/N$. Then the following assertions are equivalent:
@@
We shall show that (i) $\implies$ (ii) $\implies$ (iii) $\implies$ (i). Suppose (i) holds. Then $W=\ker(\partial)$ by Proposition 5.4. Thus $M(u)\cong\Lambda$ for every $u\in\ker(\partial)$ by Corollary 5.3 (ii) and so, in particular, assertion (ii) holds. Clearly (ii) implies (iii). Suppose (iii) holds. By Proposition 5.4, to show that (i) holds, it suffices to show that $\ker(\partial)=W$. Let $v\in\ker(\partial)$ and let $\overline{v}$ denote its image in $\ker(\partial)/N$. Then there exist (not necessarily distinct) elements $\overline{s}_{1},\ldots,\overline{s}_{r}\in S$ such that $\overline{v}=\overline{s}_{1}\cdots\overline{s}_{r}$. By hypothesis, there exist lifts $s_{1},\ldots,s_{r}\in\ker(\partial)$ such that $M(s_{1})\cong\cdots\cong M(s_{r})\cong\Lambda$, and so $s_{1},\ldots,s_{r}\in W$ by Corollary 5.3 (ii). Moreover, there exists $n\in N$ such that $v=s_{1}\cdots s_{r}n$. Since $W$ is a group and $N\leq W$, we conclude that $v\in W$. Therefore $\ker(\partial)\subseteq W$, and hence $\ker(\partial)=W$ by Corollary 5.3 (iv). ∎
5.4. A criterion in terms of sizes of locally free class groups
@@
Assume the setting of Corollary 5.3. Then there is an exact sequence
$1\longrightarrow\ker(\partial)\longrightarrow\overline{\Lambda}^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0$ (5.3)
@@
$K_{1}(\Lambda_{1})\times K_{1}(\Lambda_{2})\longrightarrow K_{1}(\overline{\Lambda})\stackrel{{\scriptstyle\partial^{\prime}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0.$ (5.4)
@@
Assume the setting of Corollary 5.3. Then we have
$|W|\leq|\ker(\partial)|=|\mathrm{Cl}(\Lambda_{1})||\mathrm{Cl}(\Lambda_{2})||\overline{\Lambda}^{\times}||\mathrm{Cl}(\Lambda)|^{-1}.$ (5.5)
@@
For any associative ring $\mathfrak{R}$ with identity, we have a short exact sequence
@@
where $J(\mathfrak{R})$ denotes the Jacobson radical of $\mathfrak{R}$ (see [CR81, Exercise 5.2], for example). In particular, if $\mathfrak{R}$ is finite then |×$|\mathfrak{R}^{\times}|=|J(\mathfrak{R})||(\mathfrak{R}/J(\mathfrak{R}))^{\times}|$. If $\mathfrak{R}$ is a finite-dimensional algebra over a finite field, then $J(\mathfrak{R})$ can be computed using the algorithm of [FR85]. We will use these observations for the computation of the cardinality of $\overline{\Lambda}^{\times}$ in certain situations.

### v3

@@ 3.1. Stably free and locally free cancellation
(3.1) $X\oplus\Lambda^{\oplus m}\cong\Lambda^{\oplus(m+n)}\textrm{ for some }m,n\geq 0\implies X\cong\Lambda^{\oplus n}.$
@@
(3.2) $X\oplus\Lambda^{\oplus m}\cong Y\oplus\Lambda^{\oplus m}\textrm{ for some }m\geq 0\implies X\cong Y.$
@@
If $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{\oplus m}\cong\Lambda^{\oplus(m+n)}\textrm{ for some }m,n\geq 0$ then it is straightforward to see that $X$ is a $\Lambda$-lattice; since $X$ is stably free it is also locally free (see §2.4). Therefore if $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders with SFC but without LFC (see [Sme15] or [SV19]).
By (2.2), $\Lambda$ has SFC if and only if (3.1) holds for all $X\in g(\Lambda)$. Similarly, $\Lambda$ has LFC if and only if (3.2) holds for all $X,Y\in g(\Lambda)$. Hence the canonical surjective map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is bijective precisely when $\Lambda$ has LFC. Moreover, $\Lambda$ has SFC precisely when the fiber above 0 has exactly one element, that is, when every stably free $\Lambda$-lattice of rank $1$ is in fact free (of rank $1$).
@@
The claim for LFC was proven by Fröhlich [Frö75, VIII], whose approach was subsequently refined by Swan. Here we adapt the proof of [Nic21a, Theorem 1.1].
By [Swa83, Theorem A10], the diagram
@@
is a weak pull back in the sense that a compatible pair of elements in $\mathrm{Cl}(\Lambda)$ and $\mathrm{LF}_{1}(\Gamma)$ has a not-necessarily-unique lift to $\mathrm{LF}_{1}(\Lambda)$. It follows that the fiber in $\mathrm{LF}_{1}(\Lambda)$ over an element $x\in\mathrm{Cl}(\Lambda)$ maps onto the fiber in $\mathrm{LF}_{1}(\Gamma)$ over the image of $x$ in $\mathrm{Cl}(\Gamma)$. Thus if $\Lambda$ has LFC then $\Gamma$ has LFC. Taking $x=0$ shows that if $\Lambda$ has SFC then $\Gamma$ has SFC. ∎
@@
Let $\Lambda_{1}$ and $\Lambda_{2}$ be $R$-orders and let $\Lambda=\Lambda_{1}\oplus\Lambda_{2}$. Then $\Lambda$ has SFC (resp. LFC) if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC (resp. LFC).
@@ 3.2. Totally definite quaternion algebras and the Eichler condition
Let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. A totally definite quaternion algebra is a simple finite-dimensional $K$-algebra $B$ whose completion at every archimedean prime of the center of $B$ is isomorphic to $\mathbb{H}$ (in particular, this forces $K$ to be totally real).
We say that a finite-dimensional semisimple $K$-algebra AA satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra. A proof of the following result can be found in [CR87, (51.24)] or [Jac68, §4].
@@
For the rest of this section, we specialize to the case of orders over the ring of integers $\mathcal{O}=\mathcal{O}_{K}$ of a number field $K$. In [SV19], it was shown that the questions of whether a given order in a totally definite quaternion algebra has SFC or LFC can be tackled algorithmically.
@@
There exists an algorithm that given an $\mathcal{O}$-order $\Lambda$ in a totally definite quaternion algebra over $K$, determines whether $\Lambda$ has SFC (resp. LFC).
@@
There exists an algorithm that given a maximal $\mathcal{O}$-order $\mathcal{M}$ in a finite-dimensional semisimple $K$-algebra, determines whether $\mathcal{M}$ has SFC (resp. LFC).
@@
We can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $\mathcal{O}$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. Therefore the desired result now follows from Corollary 3.3, Theorem 3.4 and Theorem 3.5. ∎
@@ 4. Review of cancellation properties for integral group rings
We recall results concerning SFC and LFC for integral group rings $\mathbb{Z}[G]$, where $G$ is a finite group.
@@
The finite non-cyclic subgroups of $\mathbb{H}^{\times}$ are called the binary polyhedral groups. These are the generalized quaternion groups $Q_{4n}$ for $n\geq 2$ or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups (see [CR87, §51A] and the references therein). It is well known that a finite group $G$ has a binary polyhedral quotient precisely when the group algebra $\mathbb{Q}[G]$ does not satisfy the Eichler condition (see [Nic21a, Proposition 1.5] or [CR87, (51.3)], for example).
@@
(4.1) $Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,Q_{24},\,Q_{28},\,Q_{36},\,Q_{60},\,\tilde{T},\,\tilde{O},\,\tilde{I}$.
@@
For a finite group $G$ and a field $\mathbb{F}$, let $r_{\mathbb{F}}(G)$ denote the number of irreducible $\mathbb{F}$-representations of $G$, and let $SK_{1}(\mathbb{Z}[G])=\ker(K_{1}(\mathbb{Z}[G])\rightarrow K_{1}(\mathbb{Q}[G]))$; this is a finite abelian group by a theorem of Bass [CR87, (45.20)]. We say that $K_{1}(\mathbb{Z}[G])$ is represented by trivial units if the canonical map $\pm G\rightarrow K_{1}(\mathbb{Z}[G])$ is surjective.
@@
5. Fiber products and reduction criteria for SFC
@@
5.1. Fiber products
Suppose that there is a fiber product
(5.1) $\Lambda$$\Lambda_{1}$$\Lambda_{2}$$\overline{\Lambda}.$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$
@@ 5.1. Fiber products
(5.2) $\Lambda\cong\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\quad\textrm{ and }\quad A\cong A_{1}\oplus A_{2},$
@@
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\overline{\Lambda}$ is $R$-torsion and there is a fiber product as in (5.1) such that each of $\pi_{1},\pi_{2},g_{1},g_{2}$ is the relevant canonical surjection.
@@ 5.2. Review of results of Reiner–Ullom
Given a fiber product as in (5.1), assume that at least one of $g_{1}$ and $g_{2}$ is surjective. For $u\in\overline{\Lambda}^{\times}$, define
@@ 5.3. A reduction criterion for SFC
We henceforth assume the setting and notation of Proposition 5.2 and Corollary 5.3. The following underpins many of our results.
@@
$\Lambda$ has SFC if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and $W=\ker(\partial)$.
@@
Suppose conversely that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W=\ker(\partial)$. Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free of rank $1$. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. Then for $i=1,2$ we have $\Lambda_{i}\otimes_{\Lambda}X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda_{i})$ and so $\Lambda_{i}\otimes_{\Lambda}X\cong\Lambda_{i}$ since $\Lambda_{i}$ has SFC. Hence $X\cong M(u)$ for some $u\in\overline{\Lambda}^{\times}$ by Proposition 5.2 (iv). Since $X$ is stably free, Corollary 5.3 (iii) implies that $u\in\ker(\partial)$. Thus since $W=\ker(\partial)$, Corollary 5.3 (ii) implies that $M(u)\cong\Lambda$. Therefore $X\cong M(u)\cong\Lambda$. ∎
@@
Suppose that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W$ is a group. Let $N$ be a normal subgroup of $\ker(\partial)$ contained in $W$ and let $S$ be a generating subset of $\ker(\partial)/N$. Then the following assertions are equivalent:
@@
We shall show that (i) $\implies$ (ii) $\implies$ (iii) $\implies$ (i). Suppose (i) holds. Then $W=\ker(\partial)$ by Proposition 5.4. Thus $M(u)\cong\Lambda$ for every $u\in\ker(\partial)$ by Corollary 5.3 (ii) and so, in particular, assertion (ii) holds. Clearly (ii) implies (iii). Suppose (iii) holds. By Proposition 5.4, to show that (i) holds, it suffices to show that $\ker(\partial)=W$. Let $v\in\ker(\partial)$ and let $\overline{v}$ denote its image in $\ker(\partial)/N$. Then there exist (not necessarily distinct) elements $\overline{s}_{1},\ldots,\overline{s}_{r}\in S$ such that $\overline{v}=\overline{s}_{1}\cdots\overline{s}_{r}$. By hypothesis, there exist lifts $s_{1},\ldots,s_{r}\in\ker(\partial)$ such that $M(s_{1})\cong\cdots\cong M(s_{r})\cong\Lambda$, and so $s_{1},\ldots,s_{r}\in W$ by Corollary 5.3 (ii). Moreover, there exists $n\in N$ such that $v=s_{1}\cdots s_{r}n$. Since $W$ is a group and $N\leq W$, we conclude that $v\in W$. Therefore $\ker(\partial)\subseteq W$. Since the reverse inclusion holds by Corollary 5.3 (iv), we conclude that $\ker(\partial)=W$. ∎
5.4. A criterion in terms of orders of locally free class groups
@@
There is an exact sequence
(5.3) $1\longrightarrow\ker(\partial)\longrightarrow\overline{\Lambda}^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0.$
@@
(5.4) $K_{1}(\Lambda_{1})\times K_{1}(\Lambda_{2})\longrightarrow K_{1}(\overline{\Lambda})\stackrel{{\scriptstyle\partial^{\prime}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0.$
@@
We have
(5.5) $|W|\leq|\ker(\partial)|=|\mathrm{Cl}(\Lambda_{1})||\mathrm{Cl}(\Lambda_{2})||\overline{\Lambda}^{\times}||\mathrm{Cl}(\Lambda)|^{-1}.$
@@
For any associative unital ring $\mathfrak{R}$, we have a short exact sequence
@@
where $J(\mathfrak{R})$ denotes the Jacobson radical of $\mathfrak{R}$ (see [CR81, Exercise 5.2], for example). In particular, if $\mathfrak{R}$ is finite then $|\mathfrak{R}^{\times}|=|J(\mathfrak{R})||(\mathfrak{R}/J(\mathfrak{R}))^{\times}|$. If $\mathfrak{R}$ is a finite-dimensional algebra over a finite field, then $J(\mathfrak{R})$ can be computed using the algorithm of [FR85]. We will use these observations for the computation of the cardinality of $\overline{\Lambda}^{\times}$ in certain situations.